OpenStudy (anonymous):

1.) Vivian combined different amounts of white paint, blue paint, and green paint to make 145 milliliters (ml) of paint of a new shade. The amount of white paint was x ml. The amount of blue paint was 10 ml less than twice the amount of white paint, and the amount of green paint was half the amount of blue paint. This is represented in the equation below. x + (2x – 10) + 1/2 (2x – 10) = 145 What was the difference in the amounts of white paint and blue paint Vivian combined? 40 ml 45 ml 30 ml 35 ml 2.) What is the equation of the line that passes through the points (-2, 1) and (1, 10)? 3x - y = 5 x + 3y = -5 3x - y = -5 3x - y = -7

OpenStudy (anonymous):

OH DOCTOR WHO

OpenStudy (anonymous):

Good stuff Time Lord from Gallifrey

OpenStudy (anonymous):

x + (2x-10) + 1/2(2x-10)=145 x+(2x-10)+(1x-5)=145 x+(3x-15)=145 4x-15=145 4x=160 x=40 plug into the original equation and you get 40 ml white, 70 mL blue, and 35ml green Difference from 70 to 40 is 30mL which is your answer. #2 (y^2-y^1)/(x^2-X^1) = slope (10-1)/(1--2)= 9/3 = 3 plug into y=mx+b so y=3x+b plug in one of your coordinates 10=3(1) + b 7 = b so y = 3x +7 and when you rewrite 3x - y = -7 into y = form, you get y=3x+7

OpenStudy (anonymous):

Thanks to you both XD I'm glad someone could answer my question, and know what I am :P

OpenStudy (anonymous):

NP

OpenStudy (anonymous):

Good stuff, but remember I could take you down with 1 word, no not 1, 6, 6 words, 6 words ruin you. *whispers* "Don't you think he looks tired?"

OpenStudy (anonymous):

Hmm... nope, still standing, and not ruined. :P By the way, my name is The Shadow. Nice to meet you, Mr. Kole.

OpenStudy (anonymous):

I generally go by Dr. but what sonics your screwdriver

OpenStudy (anonymous):

XD Okay then, Dr.